OpenStudy (anonymous):

let f(x)=x^4 + ax^2, what is a if f has a local min at x=5? what is the value of a if f has a point of inflection at x=2?

OpenStudy (anonymous):

\[\frac{ dy }{ dx }=4x^3+2ax\] \[\frac{ d^2y }{ dx^2 }=12x^2+2a\] At S.P's (Stationary points) \[\frac{ dy }{ dx }=0\] \[4x^3+2ax=0\] \[2x(2x^2+a)=0\] \[x=0\] or \[x^2=-\frac{a}{2}\] When y''>0, T.P(Turning Point) is minimum. \[12x^2+2a>0\] \[12(5)^2+2a>0\] \[300+2a>0\] \[2a>-300\] \[a>-150\] At I.P's (Inflexion/Inflection Points) y''=0 \[12x^2+2a=0\] \[6x^2+2a=0\] \[6(2)^2+2a=0\] \[2a=-24\] \[a=-12\]

OpenStudy (anonymous):

@Goaliess1

OpenStudy (anonymous):

@Goaliess1 You there?

OpenStudy (anonymous):

yea, I re posted it to see how others went about solving it , thanks by the way.

OpenStudy (anonymous):

Never repost.

OpenStudy (anonymous):

No matter what reason, do not repost.

OpenStudy (anonymous):

There's a bump button for a reason.

OpenStudy (anonymous):

oh ok. won't happen again.